# Each of those eight coalition logics is also determined by six other kinds of models

### v1

@@ 6.1 Clear grand-coalition-first action models
Assume (3). Assume (1) does not hold. Then, there is $a\in\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{a},\sigma^{\prime}_{a}\in\mathtt{JA}_{a}$ such that $\sigma_{a}\neq\sigma^{\prime}_{a}$ and $\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})\neq\emptyset$.
Let $w\in\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})$. Then, there is $\sigma_{\mathtt{AG}}$ and $\sigma^{\prime}_{\mathtt{AG}}$ such that $\sigma_{a}\subseteq\sigma_{\mathtt{AG}}$, $\sigma^{\prime}_{a}\subseteq\sigma^{\prime}_{\mathtt{AG}}$, $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$ and $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})$. Then, $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})\neq\emptyset$. There is a contradiction.
@@
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a clear grand-coalition-first action model, $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $\mathtt{suc}$ be the successor function of $\mathtt{G}\textbf{-}\mathtt{AM}$.
To show $\mathtt{G}\textbf{-}\mathtt{AM}$ is essentially a single-coalition-first action model, we need to show:
@@
    for all $a\in\mathtt{AG}$, for all $s\in\mathtt{ST}$, $\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$ is a general cover of $\mathtt{suc}(s)$.
@@
    for every $\mathrm{C}\subseteq\mathtt{AG}$, for all $s\in\mathtt{ST}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$,
@@
By Fact 5, (1) holds. Clearly, (3) holds. It remains to show (2).
Let $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$.
Assume $\mathrm{C}=\emptyset$. By the definition of $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})$, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\mathtt{suc}(s)$;
Assume $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$ and $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$.
By Fact 4, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\subseteq\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cup\dots\cup\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$.
Let $t\in\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cup\dots\cup\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$. Then, there is $\sigma^{1}_{\mathtt{AG}}$ such that $\sigma_{a_{1}}\subseteq\sigma^{1}_{\mathtt{AG}}$ and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{1}_{\mathtt{AG}})$, …, there is $\sigma^{n}_{\mathtt{AG}}$ such that $\sigma_{a_{n}}\subseteq\sigma^{n}_{\mathtt{AG}}$ and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{n}_{\mathtt{AG}})$. As $\mathtt{G}\textbf{-}\mathtt{AM}$ is clear, $\sigma^{1}_{\mathtt{AG}}=\dots=\sigma^{n}_{\mathtt{AG}}$. Then $\sigma_{\mathrm{C}}\subseteq\sigma^{1}_{\mathtt{AG}}$. Then $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$.
It is easy to find a single-coalition-first action model with a $\sigma_{a},\sigma^{\prime}_{a}\in\mathtt{JA}_{a}$, $\sigma_{a}\neq\sigma^{\prime}_{a}$, and $\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})\neq\emptyset$. Then it is not a clear grand-coalition-first action model.
@@
Goranko and Jamroga [4] and Ågotnes and Alechina [12] discussed injective concurrent game models, which are clear grand-coalition-first action $\mathtt{SID}$-models. As observed by Ågotnes and Alechina, the assumption behind injective concurrent game models, different available action profiles have different outcomes, is very common in game theory.
@@ 6.2 Clear single-coalition-first neighborhood models
We say $\mathtt{S}\textbf{-}\mathtt{NM}$ is clear if for all $a\in\mathtt{AG}$, and $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a partition of $\mathtt{suc}(s)$.
@@
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and let $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
@@
    for all $a\in\mathtt{AG}$, for all $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a partition of $\mathtt{suc}(s)$;
@@
    for all $\mathrm{C}\subseteq\mathtt{AG}$, for all $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathrm{C}}(s)$ is a partition of $\mathtt{suc}(s)$.
@@
Assume (1). Let $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$.
By Fact 6, $\mathtt{nei}_{\mathrm{C}}(s)$ is a general cover of $\mathtt{suc}(s)$. Assume $\mathtt{nei}_{\mathrm{C}}(s)$ is not a partition of $\mathtt{suc}(s)$. Then, there is $\mathrm{Y}_{1},\mathrm{Y}_{2}\in\mathtt{nei}_{\mathrm{C}}(s)$ such that $\mathrm{Y}_{1}\neq\mathrm{Y}_{2}$ and $\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset$. Without loss of any generality, let $w\in\mathrm{Y}_{1}$ but $w\notin\mathrm{Y}_{2}$. Let $u\in\mathrm{Y}_{1}\cap\mathrm{Y}_{2}$.
Note $\mathtt{nei}_{\emptyset}(s)=\{\mathtt{suc}(s)\}$. Then $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$.
Then, there is $X_{a_{1}}\in\mathtt{nei}_{a_{1}}(s),\dots,X_{a_{n}}\in\mathtt{nei}_{a_{n}}(s)$ such that $\mathrm{Y}_{1}=X_{a_{1}}\cap\dots\cap X_{a_{n}}$, and there is $X^{\prime}_{a_{1}}\in\mathtt{nei}_{a_{1}}(s),\dots,X^{\prime}_{a_{n}}\in\mathtt{nei}_{a_{n}}(s)$ such that $\mathrm{Y}_{2}=X^{\prime}_{a_{1}}\cap\dots\cap X^{\prime}_{a_{n}}$.
Then $w\notin X^{\prime}_{a_{i}}$ for some $1\leq i\leq n$. Note $w\in X_{a_{i}}$. Then $X_{a_{i}}\neq X^{\prime}_{a_{i}}$. Note $u\in X_{a_{i}}$ and $u\in X^{\prime}_{a_{i}}$. Then $X_{a_{i}}\cap X^{\prime}_{a_{i}}\neq\emptyset$. We have a contradiction.
@@
By 9, clear grand-coalition-first action models can be equivalently defined by three different ways, which respectively concern all agents, all coalitions and the grand coalition. It would be expected that the following condition is equivalent to the three conditions given in Theorem 11: for all $\mathtt{AG}\subseteq\mathtt{AG}$, for all $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathtt{AG}}(s)$ is a partition of $\mathtt{suc}(s)$. This is actually not the case. What follows is a counter-example.
Assume $\mathtt{AG}=\{a,b\}$. Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, where:
@@
Alurz, Henzinger, and Kupferman [2] proposed models for $\mathsf{ATL}$ based on so-called lock-step synchronous alternating transition systems. It can be shown that those models are clear single-coalition-first neighborhood $\mathtt{SID}$-models.
@@
Let $\mathtt{G}\textbf{-}\mathtt{AM}$ be a clear grand-coalition-first action model. By Theorem 10, $\mathtt{G}\textbf{-}\mathtt{AM}$ is a single-coalition-first action model. It is easy to show that $\mathtt{G}\textbf{-}\mathtt{AM}$ is $z$-representable by a single-coalition-first neighborhood model. ∎
@@
Every clear grand-coalition-first neighborhood model $z$-represents a clear single-coalition-first action model.
@@
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a clear single-coalition-first neighborhood model.
Define a single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ as in the proof for Theorem 8. As shown in that proof, $\mathtt{S}\textbf{-}\mathtt{NM}$ $z$-represents $\mathtt{S}\textbf{-}\mathtt{AM}$.
It is easy to check that $\mathtt{S}\textbf{-}\mathtt{AM}$ is clear. By Theorem 4, it is a grand-coalition-first action model.
From Theorem 12, Theorem 13, and Theorem 7, the following result follows:
Theorem 14 (Representation of clear grand-coalition-first action models by clear single-coalition-first neighborhood models).
@@
7 Tree-like grand-coalition-first action models and tree-like single-coalition-first neigh-borhood models

### v2

@@ 5.1 Clear grand-coalition-first action models
Assume (3). Suppose (1) does not hold. Then, there is $a\in\mathtt{AG}$ and $\sigma_{a},\sigma^{\prime}_{a}\in\mathtt{JA}_{a}$ such that $\sigma_{a}\neq\sigma^{\prime}_{a}$ and $\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})\neq\emptyset$. Let $t\in\mathsf{out}_{a}(s,\sigma_{a})\cap\mathsf{out}_{a}(s,\sigma^{\prime}_{a})$. Then, there is $\sigma_{\mathtt{AG}}$ and $\sigma^{\prime}_{\mathtt{AG}}$ such that $\sigma_{a}\subseteq\sigma_{\mathtt{AG}}$, $\sigma^{\prime}_{a}\subseteq\sigma^{\prime}_{\mathtt{AG}}$, $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$, and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})$. Then, $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})\neq\emptyset$, which is impossible.
For any set $\mathrm{X}$ of states, we say $\Delta\subseteq\mathcal{P}(\mathrm{X})$ is a general partition of $\mathrm{X}$ if (1) $\bigcup\Delta=\mathrm{X}$ and (2) for every $\mathrm{Y},\mathrm{Y}^{\prime}\in\Delta$, if $\mathrm{Y}\neq\mathrm{Y}^{\prime}$, then $\mathrm{Y}\cap\mathrm{Y}^{\prime}=\emptyset$. Note a general partition $\Delta$ is a partition if $\emptyset\notin\Delta$.
Fact 5.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $\mathtt{suc}$ be the sucessor function of $\mathtt{G}\textbf{-}\mathtt{AM}$.
Then, if $\mathtt{G}\textbf{-}\mathtt{AM}$ is clear, then for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$ is a general partition of $\mathtt{suc}(s)$.
Proof.
Assume $\mathtt{G}\textbf{-}\mathtt{AM}$ is clear. Let $\mathrm{C}\subseteq\mathtt{AG}$ and $t\in\mathtt{ST}$. Note $\mathtt{suc}(s)=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\mid\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}\}$. It is easy to show $\bigcup\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}=\mathtt{suc}(s)$. Let $\mathrm{X}_{1},\mathrm{X}_{2}\in\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$ such that $\mathrm{X}_{1}\neq\mathrm{X}_{2}$. Then, $\mathrm{X}_{1}=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}}^{1})$ for some $\sigma_{\mathrm{C}}^{1}\in\mathtt{JA}_{\mathrm{C}}$ and $\mathrm{X}_{2}=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}}^{2})$ for some $\sigma_{\mathrm{C}}^{2}\in\mathtt{JA}_{\mathrm{C}}$. Note $\sigma^{1}_{\mathrm{C}}\neq\sigma^{2}_{\mathrm{C}}$. Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})=\emptyset$, that is, $\mathrm{X}_{1}\cap\mathrm{X}_{2}=\emptyset$.
Note the other direction of the statement of this fact might not hold. Suppose $\mathtt{AG}=\{a\}$. Consider a grand-coalition-first action model $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ such that:
  • $\mathtt{ST}=\{s\}$;
  • $\mathtt{AC}=\{\alpha_{1},\alpha_{2}\}$;
  • $\mathsf{out}_{a}(s,\alpha_{1})=\mathsf{out}_{a}(s,\alpha_{2})=\mathtt{ST}$.
It is easy to check for all $\mathrm{C}\subseteq\mathtt{AG}$, $\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$ is a general partition of $\mathtt{suc}(s)$. However, $\mathtt{G}\textbf{-}\mathtt{AM}$ is not clear.
@@
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a clear grand-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $\mathtt{suc}$ be the successor function of $\mathtt{G}\textbf{-}\mathtt{AM}$.
To show $\mathtt{G}\textbf{-}\mathtt{AM}$ is a single-coalition-first action model, we need to show:
@@
    for all $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, $\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$ is a general cover of $\mathtt{suc}(s)$.
@@
    for all $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$,
@@
By Fact 2, (1) holds. Clearly, (3) holds. It remains to show (2). Let $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$, and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$.
Assume $\mathrm{C}=\emptyset$. Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\mathsf{out}_{\emptyset}(s,\emptyset)=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\mid\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\emptyset\subseteq\sigma_{\mathtt{AG}}\}=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\mid\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}\}=\mathtt{suc}(s)$;
Assume $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$ and $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$. By Fact 3, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\subseteq\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$. Let $t\in\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$. Then, there is $\sigma^{1}_{\mathtt{AG}}$ such that $\sigma_{a_{1}}\subseteq\sigma^{1}_{\mathtt{AG}}$ and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{1}_{\mathtt{AG}})$, …, there is $\sigma^{n}_{\mathtt{AG}}$ such that $\sigma_{a_{n}}\subseteq\sigma^{n}_{\mathtt{AG}}$ and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma^{n}_{\mathtt{AG}})$. As $\mathtt{G}\textbf{-}\mathtt{AM}$ is clear, $\sigma^{1}_{\mathtt{AG}}=\dots=\sigma^{n}_{\mathtt{AG}}$. Then $\sigma_{\mathrm{C}}\subseteq\sigma^{1}_{\mathtt{AG}}$. Then $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$.
It is easy to find a single-coalition-first action model that is not clear.
@@
Goranko and Jamroga [4] and Ågotnes and Alechina [12] discussed injective concurrent game models, which are clear grand-coalition-first action $\mathtt{SID}$-models.
As observed by Ågotnes and Alechina [12], the assumption behind injective concurrent game models, different available action profiles have different outcomes, is very common in game theory.
@@ 5.2 Clear single-coalition-first neighborhood models
We say $\mathtt{S}\textbf{-}\mathtt{NM}$ is clear if for all $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a partition of $\mathtt{suc}(s)$.
The following result gives an alternative definition of clear single-coalition-first neighborhood models.
@@
    for all $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a partition of $\mathtt{suc}(s)$;
@@
    for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathrm{C}}(s)$ is a partition of $\mathtt{suc}(s)$.
@@
Assume (1). Let $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$. By Fact 4, $\mathtt{nei}_{\mathrm{C}}(s)$ is a cover of $\mathtt{suc}(s)$. Assume $\mathtt{nei}_{\mathrm{C}}(s)$ is not a partition of $\mathtt{suc}(s)$. Then, there is $\mathrm{Y}_{1},\mathrm{Y}_{2}\in\mathtt{nei}_{\mathrm{C}}(s)$ such that $\mathrm{Y}_{1}\neq\mathrm{Y}_{2}$ and $\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset$. Without loss of any generality, we can assume there is $w$ such that $w\in\mathrm{Y}_{1}$ but $w\notin\mathrm{Y}_{2}$. Let $u\in\mathrm{Y}_{1}\cap\mathrm{Y}_{2}$. It is easy to see $\mathtt{suc}(s)\neq\emptyset$. Then, $\mathtt{nei}_{\emptyset}(s)=\{\mathtt{suc}(s)\}$, which is a singleton. Note $\mathtt{nei}_{\mathrm{C}}(s)$ is not a singleton. Then, $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$. Then, there is $\mathrm{X}_{a_{1}}\in\mathtt{nei}_{a_{1}}(s),\dots,\mathrm{X}_{a_{n}}\in\mathtt{nei}_{a_{n}}(s)$ such that $\mathrm{Y}_{1}=\mathrm{X}_{a_{1}}\cap\dots\cap\mathrm{X}_{a_{n}}$, and there is $\mathrm{X}^{\prime}_{a_{1}}\in\mathtt{nei}_{a_{1}}(s),\dots,\mathrm{X}^{\prime}_{a_{n}}\in\mathtt{nei}_{a_{n}}(s)$ such that $\mathrm{Y}_{2}=\mathrm{X}^{\prime}_{a_{1}}\cap\dots\cap\mathrm{X}^{\prime}_{a_{n}}$. Note $w\notin\mathrm{Y}_{2}$. Then $w\notin\mathrm{X}^{\prime}_{a_{i}}$ for some $a_{i}$. Note $w\in\mathrm{X}_{a_{i}}$. Then $\mathrm{X}_{a_{i}}\neq\mathrm{X}^{\prime}_{a_{i}}$. Note $u\in\mathrm{X}_{a_{i}}$ and $u\in\mathrm{X}^{\prime}_{a_{i}}$. Then $\mathrm{X}_{a_{i}}\cap\mathrm{X}^{\prime}_{a_{i}}\neq\emptyset$. We have a contradiction.
@@
By Theorem 9, clear grand-coalition-first action models can be equivalently defined in three different ways, which respectively concern all agents, all coalitions, and the grand coalition. It might be expected that the following condition, which concerns the grand coalition, is equivalent to the two conditions given in Theorem 11: for all $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathtt{AG}}(s)$ is a partition of $\mathtt{suc}(s)$. Actually, this is not the case. What follows is a counterexample.
Assume $\mathtt{AG}=\{a,b\}$. Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model such that:
@@
Alurz, Henzinger, and Kupferman [2] proposed models for $\mathsf{ATL}$ based on so-called lock-step synchronous alternating transition systems. It can be shown that they are clear single-coalition-first neighborhood $\mathtt{SID}$-models.
@@
Let $\mathtt{G}\textbf{-}\mathtt{AM}$ be a clear grand-coalition-first action model. By Theorem 10, $\mathtt{G}\textbf{-}\mathtt{AM}$ is a single-coalition-first action model. It is easy to show that $\mathtt{G}\textbf{-}\mathtt{AM}$ is $z$-representable by a clear single-coalition-first neighborhood model. ∎
@@
Every clear single-coalition-first neighborhood model $z$-represents a clear grand-coalition-first action model.
@@
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a clear single-coalition-first neighborhood model. Define a single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ as in the proof for Theorem 6. As shown in that proof, $\mathtt{S}\textbf{-}\mathtt{NM}$ $z$-represents $\mathtt{S}\textbf{-}\mathtt{AM}$. It is easy to check that $\mathtt{S}\textbf{-}\mathtt{AM}$ is clear. By Theorem 4, it is a grand-coalition-first action model.
Theorem 14 (Representation of the class of clear grand-coalition-first action $\mathtt{X}$-models by the class of clear single-coalition-first neighborhood $\mathtt{X}$-models).
@@
Proof.
Let $\mathtt{X}\in\mathrm{ES}$. Let $\mathtt{G}\textbf{-}\mathtt{AM}$ be a clear grand-coalition-first action $\mathtt{X}$-model. By Theorem 12, $\mathtt{G}\textbf{-}\mathtt{AM}$ is $z$-representable by a clear single-coalition-first neighborhood model $\mathtt{S}\textbf{-}\mathtt{NM}$. By Theorem 10, $\mathtt{G}\textbf{-}\mathtt{AM}$ is a single-coalition-first action model. By Theorem 7, $\mathtt{S}\textbf{-}\mathtt{NM}$ is a clear single-coalition-first neighborhood $\mathtt{X}$-model. Let $\mathtt{S}\textbf{-}\mathtt{NM}$ be a clear single-coalition-first neighborhood $\mathtt{X}$-model. By Theorem 13, $\mathtt{S}\textbf{-}\mathtt{NM}$ $z$-represents a clear grand-coalition-first action model $\mathtt{G}\textbf{-}\mathtt{AM}$. By Theorem 10, $\mathtt{G}\textbf{-}\mathtt{AM}$ is a single-coalition-first action model. By Theorem 7, $\mathtt{S}\textbf{-}\mathtt{AM}$ is a clear grand-coalition-first action $\mathtt{X}$-model.
6 Tree-like grand-coalition-first action models and tree-like single-coalition-first neighborhood models